# Asynchronous-Many-Task Systems: Challenges and Opportunities - Scaling an AMR Astrophysics Code on Exascale machines using Kokkos and HPX

### v1

@@ 5.5 Application of our Solutions within Octo-Tiger
Our Octo-Tiger application code was written such that it benefits from all aforementioned innovations. It was built from the ground using HPX, thus directly benefits from asynchronous task dependency construction and scheduling as well as intrinsic overlapping of computation and communication embedded in HPX’ programming model. Each grid cell of its 3D computational mesh forms a tree-node in the spanned octtree and is represented by one HPX component that can be placed on any compute node. Using HPX’s unified syntax for local and remote function calls for these components, we implemented efficient, distributed tree-traversals with little effort.
@@
While Octo-Tiger was originally developed for CPU supercomputers, we gradually ported its computational hotspots to GPUs. This required some major refactoring of the code (requiring changes to some data-structures, like moving to Struct-Of-Arrays for some of them, or moving from an interaction list to a stencil apporach for others). Given our small core developer team, this effort was done by a single developer and eventually turned Octo-Tiger it into an GPU-accelerated application [35, 1, 33, 2]. Our GPU kernels are written using Kokkos and the HPX-Kokkos integration, ensuring portability. For high efficiency on CPU platforms, we utilize C++ SIMD types within those Kokkos kernels. For instance, we use our SVE types on A64Fx systems or scalar types when compiling for GPU platforms. Thanks to the HPX-Kokkos integration, we are able to integrate the asynchronous Kokkos kernels into the HPX task-graph. This enables us to easily handle dozens of concurrent kernel launches per HPX worker thread, as we do not need to keep track of their status ourselves and let the HPX runtime handle this instead, also triggering the subsequent tasks once a Kokkos kernel is done. This automatically gains us the ability to interleave GPU kernels, CPU tasks, CPU-GPU memory transfers, and inter-node communication using the HPX task-graph. In our experience, the resulting automatic interleaving and overlapping of all of these tasks is key for achieving scalability at runtime. Notably for an GPU-accelerated application this is done without ever calling Kokkos::fence (or any equivalent synchronize method within CUDA); instead, we are relying on the event polling done by the HPX scheduling system. On GPU platforms, we implement dynamic adaptive kernel fusion (aggregation) techniques to merge individual, concurrent kernel launches into a single GPU kernel to address the GPU starvation issues. For better memory utilization, we use the recycling allocators as described above.
@@ 6.1 Tools and Utilized Scenario
To make our performance results more comparable with other codes, we include an approximation of the FLOP/s that we obtain during the runs. For this approximation, we obtain the total number of FLOPs for the scenario as follows: Table 2 shows the average floating point operations per time step over ten timesteps measured with the tool perf on an Intel Skylake CPU. We compiled Octo-Tiger (using Spack) with no vectorization support and used double precision. These obtained FLOP/s are the basis for all runs on all platforms. The meshes of level 10 and the initial mesh of level 11 fit in the memory of a single node. All other meshes required distributed runs and we measured the FLOP on a single node and multiplied it with the number of nodes. This does not accommodate the unbalanced distribution of work, however, it is a good approximation. For each node count, we executed a run for each level on all the architectures.
@@
System HPCG (PFLOP/s) Frac of Peak %
@@ 6.2 Systems and Environments
The supercomputer Fugaku is a massively parallel computer system with $158,976$ nodes based on A64FX [38] cores. The A64FX is based on the Armv8.2 architecture with scalable vector extensions (SVE). Each processor has four groups of cores called Core Memory Group (CMG) connected to a ring-bus network. Each CMG contains 12 computation cores and 1 or 0 assistant cores. A computation core has 64KiB L1I and 64KiB L1D caches. The cores in a CMG share an 8MiB L2 cache and 8GiB HMB2 memory. The total size of the memory in a node is 32GiB and the aggregate memory bandwidth is 1024GB/s. The network of Fugaku is called Tofu Interconnect D (Tofu-D), which is a hybrid 6D mesh/torus network of 40.8 GB/s total injection bandwidth per node [39]. Ookami is a small NSF-funded research cluster at Stony Brook University with 174 nodes containing A64FX CPUs and $32$ GB RAM. All nodes are connected with InfiniBand HDR 200. Except for the network architecture, it is similar to the supercomputer Fugaku. Darwin is a research testbed cluster funded by the Computational Systems and Software Environments (CSSE) subprogram of LANL’s ASC program (NNSA/DOE). It is a very heterogeneous cluster with a wide variety of hardware available, including x86, Power PC, and ARM CPU architectures, systems with terabytes of memory, and a variety of GPUs and other accelerators. For this work, we used only the NVIDIA Grace-Hopper partition of the machine. The MILK-V Pioneer is the first workstation-grade development machine for exploring RISC-V. Containing a Sophon SG2042 64-core RISC-V CPU with a 2 GHz clock frequency, 128 GB DDR 4 memory, and a 1TB PCIe 3.0 SSD.
The data artifact is available on GitHub [^2] https://github.com/STEllAR-GROUP/OperationBell24.
@@ 7.1 Alternative Communication Backend
This Octo-Tiger total time was measured over 25 time step on Perlmutter ($10$ time steps on Ookami). As production run scenarios usually take tens of thousands of time steps, we show the scaling in terms of the median runtime per timestep in Figure  3 (with a slightly better parallel efficiency of 47.647.6% than for the total time). Here, it is worth highlighting that this runtime per time step (including all computations and communication for the multiple, required solver iterations) just takes about 174174ms. Here, we benefit massively from our fine interleaving of computation and communication.
The outliers in Figure 3 are worth mentioning: Here, we can see the effect of the dynamic GPU memory pools that Octo-Tiger uses. If a GPU buffer is required the system will try to draw one from these pools and only create one if none is currently available. During the first time step, there are no buffers in the pool yet, hence it is filled over time during this first time step.
@@ 7.2 Large-Scale Runs
Figure 5 shows the achieved TFLOP/s ($10^{12}$ FLOP/s) and the number of processed cells per second on Perlmutter. Note that we show the number of processed cells per second to make the runs on Perlmutter comparable with Frontier and supercomputer Fugaku. We show the LCI results here due to the better performance. On Perlmutter, we scale up to $256$ nodes with level 10 having around $3,711$ cells per GPU per time step. For level 11, we scale up to $1,475$ nodes with an average of $6,814$ cells per GPU per time step. The full system of 40 GB GPUs is $1,536$, however, $1,475$ were available during our reservation for the full system run. Combining the 40 GB GPUs and the 80GB GPUs with a higher memory bandwidth allowed for the largest run of $1,720$ nodes. Here, each GPU had on average $37,398$ cells per GPU per time step.
@@
Figure 7 shows the achieved TFLOP/s and the number of processed cells per second on supercomputer Fugaku for all three levels. We used the SVE backend for our SIMD types. Level 10 scaled from a single node up to 2048 nodes with around $40$ cells per core per time step. Level 11 fit in $8$ nodes and scaled up to $6,000$ nodes with around $140$ cells per core per time step. Level 12 runs from $1,024$ nodes up to $6,144$ nodes with around $872$ cells per core per time step. We need to mention here that going from level 11 to level 12 the number of cells increased by four and the memory usage by $6$, see Table 2. Note that the largest run was on $294,912$ cores. Here, we reached a total time $97.14$s with a parallel efficiency 64.564.5% (with respect to the smallest level 12 run using $1,024$ nodes). A notable difference is that the Fujitsu MPI used on supercomputer Fugaku only supports MPI_THREAD_SERIALIZED and we had to add additional synchronization in our code. However, we could use MPI_THREAD_MULTIPLE for the other systems. The additional synchronization overhead shows up for level 12 with many more messages. For a detailed study on the MPI threading level, we refer to [5]. We observe fewer TFLOP/s for level 12 than for level 11 on A64FX, which we do not observe on Frontier and Perlmutter. For this reason, we have not performed runs on even larger node counts on supercomputer Fugaku.
@@
We observed the lowest TFLOP/s on supercomputer Fugaku compared to the two machines with GPU acceleration. The full system run on Perlmutter using mixed GPUs had an order of magnitude higher TFLOP/s as $1,024$ Frontier nodes.
@@ 8 Conclusion and Outlook
We have created Octo-Tiger, an astrophysical code that can simulate astrophysical phenomena such as the merger of binary star systems. Exemplary for many applications, we need to employ dynamic adaptive mesh refinement to efficiently bridge across multiple physical scales. In contrast to alternative approaches based on smooth particle hydrodynamics, a mesh-based finite volume scheme allows us to conserve important physical quantities such as momentum and energy up to machine precision. Additionally, it enables convergence studies that are critical to a deeper understanding of the underlying phenomena. We realize, for the first time, highly resolved simulations with up to 257 million dynamically adaptive cells, which was previously infeasible. With that we have prepared the ground for extreme-scale simulations to gain a deeper understanding of the evolution of our universe.
@@
In our largest runs, we have achieved for $257$ million dynamically adaptive cells 51.3751.37% parallel efficiency on Frontier on $32,768$ cores and $2,048$ MI250X GPUs with about 1717% HPCG peak performance; 64.4764.47% on $294,912$ cores on supercomputer Fugaku (55% HPCG peak); and 47.5947.59% for a full system run on Perlmutter on $4,110,080$ cores, and $6,880$ A100 GPUsusing both the available 40GB A100 nodes with the 80GB A100 nodes simultaneously (2626% HPCG peak).
Our contributions have also brought astrophysical simulations into a unique position to gain code and performance portability for novel and future systems. On a node level, we already have shown code portability to the new Intel GPUs, to Raspberry PIs, and to RISC-V CPUs. These efforts are pivotal in laying the groundwork for its use on tomorrow’s upcoming ARM64 and RISC-V HPC systems such as the European RISC-V flagship supercomputer announced for 2026.
All of our computational and algorithmic achievements are not specific to our guiding astrophysical application. The insights and significant improvements gained can be directly transferred to other mesh-based multi-physics codes with dynamic adaptivity, thus promising a high impact on future code developments.
However, while the scalability and portability results with Octo-Tiger look extremely promising, there is still much to improve within the application itself, as is shown by the achieved FLOP/s: One major downside of the current version of Octo-Tiger, is that it is a GPU-accelerated application, not yet a GPU-resident one. While its computational hotspots within the hydro and gravity solvers have been ported to Kokkos (and thus to GPUs), there are many parts of its code that still run on CPU. This means that the GPU results constantly need to be transferred to the CPU and be processed there by these other code parts. Worse, as we had to refactor some utilized data-structures to make Octo-Tiger more suitable for GPUs. (moving from Array-of-Structs to Struct-of-Arrays, for example, see: [40]), the interface between these older and newer parts sometimes includes data-structure conversions. All of this puts a lot of stress on the main memory bandwidth, turning it into a bottleneck. This is mostly a result of refactoring and porting Octo-Tiger piece by piece. Despite this bottleneck, we still achieve major speedups when using GPUs over CPUs [1, 34, 2, 33], however, the aforementioned legacy CPU code parts prevent us from realizing Octo-Tiger’s full potential on GPU machines just yet.
@@
Acknowledgments
This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231 using NERSC award DDR-ERCAP0028472. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. This research used computational resources of the supercomputer Fugaku provided by RIKEN Center for Computational Science. This work was supported by the U.S. Department of Energy through the Los Alamos National Laboratory (LANL). LANL is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). We also thank the LANL Advanced Simulation & Computing Program and CCS-7 Darwin cluster for computational resources. The authors would like to thank Stony Brook Research Computing and Cyberinfrastructure, and the Institute for Advanced Computational Science at Stony Brook University for access to the innovative high-performance Ookami computing system, which was made possible by a \$5M National Science Foundation grant (#1927880). The support we received from the Center of Computation and Technology at Louisiana State University was invaluable. The authors also acknowledge the technical support we received from NVIDIA (Scot Halverson) in the early stages of the project. Assigned: LA-UR-24-23457 (Rev. 2). Notice of copyright: This manuscript has been authored by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S. Department of Energy. The United States Government retains and the publisher, by accepting the article for publication, acknowledges that the United States Government retains a non-exclusive, paid-up, irrevocable, worldwide license to publish or reproduce the published form of this manuscript, or allow others to do so, for United States Government purposes. The Department of Energy will provide public access to these results of federally sponsored research in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan).

### v2

@@ 5.5 Application of our Solutions within Octo-Tiger
Our Octo-Tiger application code was written such that it benefits from all aforementioned innovations. It was built from the ground using HPX, thus directly benefits from asynchronous task dependency construction and scheduling as well as intrinsic overlapping of computation and communication embedded in HPX’s programming model. Each grid cell of its 3D computational mesh forms a tree-node in the spanned octtree and is represented by one HPX component that can be placed on any compute node. Using HPX’s unified syntax for local and remote function calls for these components, we implemented efficient, distributed tree-traversals with little effort.
@@
While Octo-Tiger was originally developed for CPU-based supercomputers, we gradually ported its computational hotspots to GPUs. This required some major refactoring of the code (requiring changes to some data-structures, like moving to Struct-Of-Arrays for some of them, or moving from an interaction list to a stencil approach for others). Given our small core developer team, this effort was done by a single developer and eventually turned Octo-Tiger into a GPU-accelerated application. However, it is not yet a GPU-resident code: Only the computational-intensive kernels were ported to Kokkos and thus to GPUs, the rest of the application still runs on the CPU. While this already helps us to achieve major GPU speedups with Octo-Tiger [43, 1, 41, 2], it also means that the data is still primarily located on the main memory, not the GPU memory. Thus, there are numerous parts with CPU-GPU communication as the execution flow goes from GPU back to CPU execution, which puts more stress on the CPU-GPU connection and the main memory bandwidth itself, limiting the overall performance achieved by Octo-Tiger. Although there is more work to be done on Octo-Tiger, the existing GPU kernels already speed up the application noticeably in terms of total application runtime. In the context of this work, the frequent CPU-GPU data-transfers and the interleaving of numerous fine-grained CPU and GPU tasks actually provide a good stress test for our integrations and the HPX task-graph processing itself. Regarding Octo-Tiger’s GPU kernels themselves, the kernels are written using Kokkos and our HPX-Kokkos integration, ensuring portability. For high efficiency on CPU platforms, we utilize C++ SIMD types within those Kokkos kernels. For instance, we use our SVE types on A64FX systems or scalar types when compiling for GPU platforms. Thanks to the HPX-Kokkos integration, we are able to integrate the asynchronous Kokkos kernels into the HPX task-graph. This enables us to easily handle dozens of concurrent kernel launches per HPX worker thread, as we do not need to keep track of their status ourselves and let the HPX runtime handle this instead, also triggering the subsequent tasks once a Kokkos kernel is done. This automatically gains us the ability to interleave GPU kernels, CPU tasks, CPU-GPU memory transfers, and inter-node communication using the HPX task-graph. In our experience, the resulting automatic interleaving and overlapping of all of these tasks is key for achieving scalability at runtime. Notably, for a GPU-accelerated application, this is done without ever calling Kokkos::fence (or any equivalent synchronize method within CUDA); instead, we are relying on the event polling done by the HPX scheduling system. On GPU platforms, we implement dynamic adaptive kernel fusion (aggregation) techniques to merge individual, concurrent kernel launches into a single GPU kernel to address the GPU starvation issues. For better memory utilization, we use the recycling allocators as described above.
@@ 6.1 Tools and Utilized Scenario
To make our performance results more comparable with other codes, we include an approximation of the FLOP/s that we obtain during the runs. For this approximation, we obtain the total number of FLOPs for the scenario as follows: Table 3 shows the average floating point operations per time step over ten timesteps measured with the tool perf on an Intel Skylake CPU. We compiled Octo-Tiger (using Spack) with no vectorization support and used double precision. These obtained FLOP/s are the basis for all runs on all platforms. The meshes of level 10 and the initial mesh of level 11 fit in the memory of a single node. All other meshes required distributed runs and we measured the FLOP on a single node and multiplied it with the number of nodes. This does not accommodate the unbalanced distribution of work. However, as we distribute the workload on as few compute nodes as possible to accommodate for the memory demands, the individual workload per compute node is large enough to make this a good enough approximation for our purposes nonetheless. This approximation is not exact, but provides us with a first estimate of our achieved FLOP/s.
@@
System
HPCG
(PFLOP/s)
Frac of Peak
@@ 6.2 Systems and Environments
The supercomputer Fugaku is a massively parallel computer system with $158,976$ nodes based on A64FX [46] cores. The A64FX is based on the Armv8.2 architecture with scalable vector extensions (SVE). Each processor has four groups of cores called Core Memory Group (CMG) connected to a ring-bus network. Each CMG contains 12 computation cores and 1 or 0 assistant cores. A computation core has 64KiB L1I and 64KiB L1D caches. The cores in a CMG share an 8MiB L2 cache and 8GiB HMB2 memory. The total size of the memory in a node is 32GiB and the aggregate memory bandwidth is 1024GB/s. The network of Fugaku is called Tofu Interconnect D (Tofu-D), which is a hybrid 6D mesh/torus network of 40.8 GB/s total injection bandwidth per node [47]. Ookami is a small NSF-funded research cluster at Stony Brook University with 174 nodes containing A64FX CPUs and $32$ GB RAM. All nodes are connected with InfiniBand HDR 200. Except for the network architecture, it is similar to the supercomputer Fugaku. Darwin is a research testbed cluster funded by the Computational Systems and Software Environments (CSSE) subprogram of LANL’s ASC program (NNSA/DOE). It is a very heterogeneous cluster with a wide variety of hardware available, including x86, Power PC, and ARM CPU architectures, systems with terabytes of memory, and a variety of GPUs and other accelerators. For this work, we used only the NVIDIA Grace-Hopper partition of the machine. The MILK-V Pioneer is the first workstation-grade development machine for exploring RISC-V. Containing a Sophon SG2042 64-core RISC-V CPU with a 2 GHz clock frequency, 128 GB DDR 4 memory, and a 1TB PCIe 3.0 SSD. Table 5 summarizes the used compilers and environments. The data artifact is available on GitHub [^3] https://github.com/STEllAR-GROUP/OperationBell24.
System compiler mpi hpx kokkos hpx-kokkos GPU SDK lci
Fugaku gcc 12.2.0 fujitsu-mpi head 1.9.1 4.0.01 0.4.0
Perlmutter gcc 12.3.0 cray-mpich 8.1.28 310f449 4.0.01 0.4.0 CUDA 12.2 6dcb6a8
Frontier rocmcc 5.7.0 cray-mpich 8.1.28 310f449 4.0.01 0.4.0 HIP 5.7.0
Ookami gcc 12.2.0 openmpi 4.1.5 40eef48 4.1.00 0.4.0 6dcb6a8
Table 5: Software environments used on the systems. Some systems required patches and we provided the short git hash in italic.
@@ 7.1 Alternative Communication Backend
This Octo-Tiger total time was measured over 25 time steps on Perlmutter ($10$ time steps on Ookami). As production run scenarios usually take tens of thousands of time steps, we show the scaling in terms of the median runtime per timestep in Figure  4 (with a slightly better parallel efficiency of 47.647.6% than for the total time). Here, it is worth highlighting that this runtime per time step (including all computations and communication for the multiple, required solver iterations) just takes about 174174ms. Here, we benefit massively from our fine interleaving of computation and communication.
The outliers in Figure 4 are worth mentioning: Here, we can see the effect of the dynamic GPU memory pools that Octo-Tiger uses. If a GPU buffer is required, the system will try to draw one from these pools and only create one if none is currently available. During the first time step, there are no buffers in the pool yet, hence it is filled over time during this first time step.
@@ 7.2 Large-Scale Runs
Figure 6 shows the achieved TFLOP/s ($10^{12}$ FLOP/s) and the number of processed cells per second on Perlmutter. Note that we show the number of processed cells per second to make the runs on Perlmutter comparable with Frontier and supercomputer Fugaku. We show the LCI results here due to the better performance. On Perlmutter, we scale up to $256$ nodes with level 10 having around $3,711$ cells per GPU per time step. For level 11, we scale up to $1,475$ nodes with an average of $6,814$ cells per GPU per time step. Though Perlmutter features $1,536$ 40 GB GPUs, only $1,475$ were available during our reservation for the full system run. Combining the 40 GB GPUs and the 80GB GPUs with a higher memory bandwidth allowed for the largest run of $1,720$ nodes. Here, each GPU had on average $37,398$ cells per GPU per time step.
@@
Figure 8 shows the achieved TFLOP/s and the number of processed cells per second on supercomputer Fugaku for all three levels. We used the SVE backend for our SIMD types. Level 10 scaled from a single node up to 2048 nodes with around $40$ cells per core per time step. Level 11 fit in $8$ nodes and scaled up to $6,000$ nodes with around $140$ cells per core per time step. Level 12 runs from $1,024$ nodes up to $6,144$ nodes with around $872$ cells per core per time step. We need to mention here that going from level 11 to level 12 the number of cells and the memory usage increased by $6.4$, see Table 3. Note that the largest run was on $294,912$ cores. Here, we reached a total time $97.14$s with a parallel efficiency 64.564.5% (with respect to the smallest level 12 run using $1,024$ nodes). A notable difference is that the Fujitsu MPI used on supercomputer Fugaku only supports MPI_THREAD_SERIALIZED and we had to add additional synchronization in our code. However, we could use MPI_THREAD_MULTIPLE for the other systems. The additional synchronization overhead shows up for level 12 with many more messages. For a detailed study on the MPI threading level, we refer to [5]. We observe fewer TFLOP/s for level 12 than for level 11 on A64FX, which we do not observe on Frontier and Perlmutter. For this reason, we have not performed runs on even larger node counts on supercomputer Fugaku.
@@
We observed the lowest TFLOP/s on supercomputer Fugaku compared to the two machines with GPU acceleration. The full system run on Perlmutter using mixed GPUs achieved almost twice the number of TFLOP/s ($500.03$ TFLOP/s) as the run using $1024$ Frontier nodes ($255.29$ TFLOP/s).
@@ 8 Conclusion and Outlook
We have created Octo-Tiger, an astrophysical code that can simulate astrophysical phenomena such as the merger of binary star systems. Exemplary for many applications, we need to employ dynamic adaptive mesh refinement to efficiently bridge across multiple physical scales. In contrast to alternative approaches based on smooth particle hydrodynamics, a mesh-based finite volume scheme is used to allow us to conserve important physical quantities such as momentum and energy up to machine precision. Additionally, it enables convergence studies that are critical to a deeper understanding of the underlying phenomena. We realize, for the first time, highly resolved simulations with up to 257 million dynamically adaptive cells, which was previously infeasible. With this work, we have prepared the groundwork for extreme-scale simulations to gain a deeper understanding of the evolution of our universe.
@@
In our largest runs with $257$ million dynamically adaptive cells, we have achieved 51.3751.37% parallel efficiency on Frontier on $32,768$ cores and $2,048$ MI250X GPUs with about 1717% HPCG peak performance; 64.4764.47% on $294,912$ cores on supercomputer Fugaku (55% HPCG peak); and 47.5947.59% for a full system run on Perlmutter on $4,110,080$ cores, and $6,880$ A100 GPUs using both the available 40GB A100 nodes with the 80GB A100 nodes simultaneously (2626% HPCG peak).
Our contributions have also brought astrophysical simulations into a unique position to gain code and performance portability for novel and future systems. On a node level, we have already shown code portability to the new Intel GPUs, to Raspberry PIs, and to RISC-V CPUs. These efforts are pivotal in laying the groundwork for its use on tomorrow’s upcoming ARM64 and RISC-V HPC systems such as the European RISC-V flagship supercomputer announced for 2026.
These computational and algorithmic achievements are not specific to our guiding astrophysical application. The insights and significant improvements gained can be directly transferred to other mesh-based multi-physics codes with dynamic adaptivity, thus promising a high impact on future code developments.
However, while the scalability and portability results with Octo-Tiger look extremely promising, there is still much to improve within the application itself, as is shown by the achieved FLOP/s. One major downside of the current version of Octo-Tiger, is that it is a GPU-accelerated application, not yet a GPU-resident one. While its computational hotspots within the hydro and gravity solvers have been ported to Kokkos (and thus to GPUs), there are many parts of its code that still run on CPU. This means that the GPU results constantly need to be transferred to the CPU to be processed there by other parts of code. Worse, as we had to refactor some utilized data-structures to make Octo-Tiger more suitable for GPUs (moving from Array-of-Structs to Struct-of-Arrays, for example, see: [48]), the interface between these older and newer parts sometimes includes data-structure conversions. This puts a lot of stress on the main memory bandwidth, turning it into a bottleneck. This is mostly a result of refactoring and porting Octo-Tiger piece by piece. Despite this bottleneck, we still achieve major speedups when using GPUs over CPUs [1, 42, 2, 41]. However, the aforementioned legacy CPU code parts prevent us from realizing Octo-Tiger’s full potential on GPU machines just yet.
@@
Acknolegments
This research used resources of the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility located at Lawrence Berkeley National Laboratory, operated under Contract No. DE-AC02-05CH11231 using NERSC award DDR-ERCAP0028472. This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. This research used computational resources of the supercomputer Fugaku provided by RIKEN Center for Computational Science. This work was supported by the U.S. Department of Energy through the Los Alamos National Laboratory (LANL). LANL is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001). We also thank the LANL Advanced Simulation & Computing Program and CCS-7 Darwin cluster for computational resources. The authors would like to thank Stony Brook Research Computing and Cyberinfrastructure, and the Institute for Advanced Computational Science at Stony Brook University for access to the innovative high-performance Ookami computing system, which was made possible by a \$5M National Science Foundation grant (#1927880). The support we received from the Center of Computation and Technology at Louisiana State University was invaluable. The authors also acknowledge the technical support we received from NVIDIA (Scot Halverson) in the early stages of the project. Assigned: LA-UR-24-23457 (Rev. 3).